use the intercept form to find the equation on the line that contains the point
(-8,2) and has slope 3/4

Answer by fractalier(6550)

The slope intercept form is
y - y1 = m(x - x1)
and we get
y - 2 = (3/4)(x + 8)
y - 2 = (3/4)x + 6
y = (3/4)x + 8